OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 3x + y = 6 4x − 5y = 10

OpenStudy (anonymous):

Isolating y in the first equation and then substituting -3x+6 for y in the second function

OpenStudy (anonymous):

can u show me step by step

OpenStudy (anonymous):

3x+y=6 -3x -3x y=-3x+6 Now sustitute

OpenStudy (anonymous):

umm i dont know how

OpenStudy (anonymous):

4x-5(-3x+6)=10 distribute the 5 4x+15x-30=10 combine like terms

OpenStudy (anonymous):

4x+15x-30=10 19x=40 x=19/40

OpenStudy (anonymous):

sorry that would be x=40/19

OpenStudy (anonymous):

Does all that make sense?

OpenStudy (anonymous):

yeah tank u

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

thats not an answer

OpenStudy (anonymous):

Whats the answer then?

OpenStudy (anonymous):

You asked for the first step. The very first step is to put the first function in y=mx+b format

OpenStudy (anonymous):

You could always set the equations equal to each other

OpenStudy (anonymous):

so my first step is to multipy the equation by 5

OpenStudy (anonymous):

or to add the two equations together

OpenStudy (anonymous):

Ohh I thought your goal was to combine and solve for x. SO you are actually do addition of the functions....

OpenStudy (anonymous):

Multiply the first equation by 5. Multiply the first equation by −5. Multiply the first equation by −4. Add the two equations together.

OpenStudy (anonymous):

which one is the step lol

OpenStudy (anonymous):

So by multiplying by 5 you would be able to get rid of the y by having on -5 and one positive. Well looks like you figured out the first step :)

OpenStudy (anonymous):

so multipyln 5